C-CHART To monitor the number of blemishes on a polished surface a company randomly selects 10...

C-CHART

To monitor the number of blemishes on a polished surface a company randomly selects 10 units of output from its process and counts the number of blemishes on each unit. The sample results are shown below

SAMPLE NO. 1 2 3 4 5 6 7 8 9 10
# BLEMISHES 17 19 15 18 16 14 15 16 15 15

Determine the upper and lower control limits for 3-sigma control chart to monitor the number of blemishes per unit
Answer #1

c (True defective) = Total number of defects/Number of sample = (17+19+15+....................+15+15)/10 = 160/10 = 16

standard deviation (s) = sqrt(c) =sqrt(16) = 4

z = 3

UCL = c + z*s = 16+3*4 = 28

LCL = c - z*s = 16 - 3*4 = 4